OpenStudy (anonymous):

@iGreen help i will meadel and fan c/4-5=4

OpenStudy (igreen):

Is this it? \(\dfrac{c}{4} - 5 = 4\) Or is it: \(\dfrac{c}{4-5} = 4\)

OpenStudy (anonymous):

um it is the one above or is it

OpenStudy (igreen):

Is it the first one or 2nd one there?

OpenStudy (anonymous):

first

OpenStudy (igreen):

Okay, first step is to add 5 to both sides, what's 4 + 5?

OpenStudy (anonymous):

9

OpenStudy (igreen):

Yes, so we have: \(\dfrac{c}{4} = 9\) Now multiply 4 to both sides, what's 9 * 4? @mippy2